For the cell reaction represent as
Co/Cos+//Not+/Ni
Calculate the cell potential and free energy of reaction
1
Expert's answer
2020-02-13T06:15:36-0500

Standard electrode potential Co/Co2+E°L=-0.7 V and Ni2+/Ni E°R=-0.24

Cell potential:

"E\u00b0=E\u00b0R-E\u00b0L"

E°=-0.24-(-0.27)=0.03 (V)

Free energy or Gibbs energy:

"\u2206G=-n\u2219F\u2219E\u00b0" ,

n is the number of electrons participating in the reaction; F-Faraday number 96485 C∙mol-1

∆G=-2∙96485-0.03=-5789.1 J/mol